OpenStudy (anonymous):

Lim x to inf (1+ 1/x)^(2x) Without L'Hospital's

OpenStudy (anonymous):

what is \[\lim_{x\to \infty}(1+\frac{1}{x})^x\]?

OpenStudy (anonymous):

the power is 2x and the result is 1 i dont understand why it's not e^2 =(

OpenStudy (anonymous):

it is \(e^2\)

OpenStudy (anonymous):

oops sorry, wrong number

OpenStudy (anonymous):

no wonder you don't understand why it is not \(e^2\) since it is !!

OpenStudy (anonymous):

its lim x to 0 (1+2x)^x

OpenStudy (anonymous):

this one is 1 and I dont understand!!!

OpenStudy (anonymous):

what form is this in?

OpenStudy (anonymous):

@eliassaab where you been? nice to see you back

OpenStudy (anonymous):

Thanks, I have been busy travelling.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \left( 1+2x \right)^{x}\]

OpenStudy (anonymous):

sounds nice @matheusoliveira i mean "what form" as in \(\frac{0}{0}\) or \(1^{\infty}\) for example

OpenStudy (anonymous):

i dont know i think its 1^0 but i dont know if its right

OpenStudy (anonymous):

yes you are right and that form is not "indeterminate" that form is just 1

OpenStudy (anonymous):

i see, can u help some more? pliz